17. (x+5)(x-3) 18. (7r+9)(5r-3)
Answer #1

G a 8 a 2 аха 5 3 了 . (x+s) (z-3 = x (x-3) +5(x-3 ) 2 18..(לך1+( ) 57-3) = 7 (50-3) +1 (57-3) 3502t 247-27 (57-3)= 3504247-27